8. Evaluate # F. dS if F(x, y, z) = (2x + 24 + xy, z tanh-lz - y - y?, zy+ x +y In sy 2 and S is the boundary of the solid enplease use completely simplified exact values only, thank you.
Answer #1

Gauss Divergence Theorem It states that the surface integral of the normal component of a vector function Å taken over closedi divf = (2+ y) + (-1-2y) + (4-1) = 2+ y - 1 - 2y + y-1 $ Fids = Isso dv S >> F.ds = 0 s